What are the major sources of economic profit? Oa. certainty, monopolistic competition, and inelastic supply. Ob. competition, perfect information, and elasticity of market demand. oc. barriers to entry, uncertainty, and entrepreneurial alertness. Od. externalities, inflation, and size of firm.
The procedure used to calculate the present value of future income is called Oa. indirect production. O b. investment tax shelter planning. Oc. discounting. d. amortizing.
If the interest rate is 5 percent, the net present value of $200 to be received two years from now is Oa. $165.29 Ob. S181.41 Oc. $200. Od. S210.
Answer #1

Q1:

  • Option B is incorrect because all three will contribute to only normal profit
  • Option D is incorrect because all three will are not in the control of the supplier and lead to lack of sales which in turn leads to lack of profit
  • Option A is incorrect because all three will contribute to only normal profit

Therefore by process of elimination, option C is correct. Let's look at it logically, with barriers to entry no new players can come in, uncertainty also makes new players averse from entering the market and entrepreneurial alertness will also add to the possibilities of economic profit.

Q2: As finding the future value is done by compounding the opposite process is called discounting. Following equations represent the two processes:

Compounding: FV = CF \times (1+r)^{n}

Discounting: CF NPV = (1 +r)n

So option C is correct

Q3: To calculate the NPV we need to solve the following equation:

CF NPV = (1+r)2

200 NPV = (1.05)2

NPV-181.41

So option B is correct